OpenStudy (anonymous):

I have this problem that I am working on √5(6√2-6√25. I took what is inside of the parenthesis and then I substract

OpenStudy (anonymous):

\[\sqrt{5}\times(6\sqrt{2}-6\sqrt{25})?\]

OpenStudy (anonymous):

sorry I forgot to put the last parenthsis